Opened 7 years ago

Last modified 9 months ago

#470 new flight processing

ET12/17, flight day 321/2012, Corbetti_Alutu

Reported by: mark1 Owned by: benj
Priority: alpha 4 high Milestone:
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by anhi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/ethiopia/ET12_17-2012_321_Corbetti_Alutu/

Data arrived from ARSF via SATA disk on 06/12/2012

ET12/17 Scientific objective:

ET12/17 Priority: a4h

PI: David Pyle

Sensors:

Camera (05/03/2013)
Eagle (Requested)
Hawk (Requested)
LiDAR (05/03/2013)

Change History (38)

comment:1 Changed 7 years ago by mark1

GPS basestation data acquired via ftp 17/12/12

comment:2 Changed 7 years ago by mark1

Note to whoever starts processing this project - it may be simpler in the long run to split this project up into 2 [a for corbetti, b for alutu]. See how the navigation processing goes.

comment:3 Changed 7 years ago by tipo

Started Hyperspectral Processing.
used ppp for navigation, but could only get '2's on GrafNav

comment:4 Changed 6 years ago by tipo

Finished Hyperspectral processing.

comment:5 Changed 6 years ago by edfi

Starting Hyperspectral delivery check

comment:6 Changed 6 years ago by edfi

Digital logsheet needs to be generated and added into the delivery.
Is there supposed to be a h321191? I can't see one, meaning uneven numbers of lines between eagle and hawk. Is this correct?
All else seems fine

comment:7 Changed 6 years ago by tipo

There's a line listed in the logsheet as "HAWK FAULT" which must be why the lines are uneven.

comment:8 Changed 6 years ago by tipo

PDF logsheet generated an put in delivery.

comment:9 Changed 6 years ago by besm

Starting hyperspectral delivery check.

Last edited 6 years ago by besm (previous) (diff)

comment:10 Changed 6 years ago by besm

Delivery check completed. The progress details of the ticket show this has been delivery checked before, and there's a tmp_cmd_check directory already present, but I saw nothing written here so I went ahead and did everything. Issues found:

  • Dubious about the sct values of eagle lines 11 & 16, and hawk line 15. (I looked at the existing tmp_cmd_check directory. If this has been fixed, tell me and I will do that step again.)
  • Looks like there's a dropped scan on eagle line 8 that'd be worth mentioning in the readme.

comment:11 follow-up: Changed 6 years ago by edfi

Current basestations not currently good enough, waiting for new data

comment:12 in reply to: ↑ 11 Changed 6 years ago by edfi

Replying to edfi:

Current basestations not currently good enough, waiting for new data

(For LiDAR processing)

comment:13 Changed 6 years ago by mark1

Raw basestation files arrived (ftp) - copied into basestation/raw

comment:14 Changed 6 years ago by edfi

Re-attempting navigation with new basestation data

comment:15 Changed 6 years ago by edfi

Nav complete, default settings used in alspp. Still need to look to see if pitch/roll adjustment is required

comment:16 Changed 6 years ago by anhi

RCD Delivery Created.

comment:17 Changed 6 years ago by edfi

Starting RCD check

comment:18 Changed 6 years ago by edfi

Check complete. Ready to deliver

comment:19 Changed 6 years ago by tipo

Corrected hyperspectral lines e11, e16 and h15 and mentioned the dropped scan in the readme.

comment:20 Changed 6 years ago by benj

PI has requested that this flight be processed using the December 2012 calibration - please process the hyperspectral data accordingly.

comment:21 Changed 6 years ago by edfi

LiDAR Ready for delivery check

comment:22 Changed 6 years ago by anhi

LiDAR Delivery check:

No problems, ready for delivery.

comment:23 Changed 6 years ago by anhi

Delivered LiDAR and rcd data on 05/03/2013 on disk 152 to David Pyle, University of Oxford.

comment:24 Changed 6 years ago by anhi

  • Description modified (diff)

comment:25 Changed 6 years ago by tipo

Reprocessed Hyperspectral with new calibration.

comment:26 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • No logsheet in delivery
  • Hawk line Ul001 is missing not line Ul002, needs updating in Readme
  • Mapped files incorrectly named

comment:27 Changed 6 years ago by tipo

Logsheet added, readme updated, incorrectly named files fixed.

comment:28 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • Logsheet not filled in.
  • Quality and issues of data part of readme sill says hawk line UL002 is missing.

No other problems

comment:29 Changed 6 years ago by tipo

Corrected the readme and replaced the logsheet file with the correct one.

comment:30 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • Mapped files still have the sct in the file name.

No other problems

comment:31 Changed 6 years ago by tipo

sct removed from the filenames

comment:32 Changed 6 years ago by anhi

Hyperspectral delivery check:

  • Removed a copy of the dem
  • Finished renaming and zipped the mapped files

No other problems found, ready for delivery

comment:33 Changed 6 years ago by tipo

Delivered hyperspectral data to David Pyle, University of Oxford 3/5/2013

comment:34 Changed 6 years ago by tipo

archiving

comment:35 Changed 6 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Owner set to benj

comment:36 Changed 6 years ago by tipo

Renamed /posatt/ipas_honeywell/extract/2012321_RTMsg.txt to 2012_321_RTMsg.txt as requested by proj_tidy.sh

comment:37 Changed 5 years ago by tipo

Made avalable for archiving 13/01/2014

comment:38 Changed 9 months ago by dac

Archiving

Data available from CEDA. Closing ticket.

Note: See TracTickets for help on using tickets.